OpenStudy (anonymous):

Here is another radical equation, this one is a bit different than the others

OpenStudy (anonymous):

\[\sqrt{-3x+16}-1=\sqrt{-2x+9}\]

OpenStudy (anonymous):

I'm confused about the 1 in the middle mostly

OpenStudy (vijay):

x^ 2 - 10x + 32 = 0 u will get this then solve it

OpenStudy (anonymous):

so do i factor it?

OpenStudy (vijay):

1st square both side of the equation,

OpenStudy (vijay):

-3x + 16 + 1 - 2( -3x +16)^0.5 = -2x + 9 -x + 8 -2 (-3x +16)^0.5 = 0 8-x = 2 (-3x +16)^0.5 square both side again 64 - 16x + x^2 = 2(-3x+16) = -6x +32 32 -10 + X^2 = 0 factorise this